# A Roadmap for Software Testing in Open Collaborative Development Environments

### v1

@@ Abstract.
Amidst the ever-expanding digital sphere, the evolution of the Internet has not only fostered an atmosphere of information transparency and sharing but has also sparked a revolution in software development practices. The distributed nature of open collaborative development, along with its diverse contributors and rapid iterations, presents new challenges for ensuring software quality. This paper offers a comprehensive review and analysis of recent advancements in software quality assurance within open collaborative development environments. Our examination covers various aspects, including process management, personnel dynamics, and technological advancements, providing valuable insights into effective approaches for maintaining software quality in such collaborative settings. Furthermore, we delve into the challenges and opportunities arising from emerging technologies such as LLMs and the AI model-centric development paradigm. By addressing these topics, our study contributes to a deeper understanding of software quality assurance in open collaborative environments and lays the groundwork for future exploration and innovation.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: International Workshop on Software Engineering in 2030; November 2024; Puerto Galinàs (Brazil)
@@ 1. Introduction
Open collaboration is a hallmark of the Internet era in the past ten to twenty years, where geographical distances cease to be a barrier. Platforms like social media networks and online forums have emerged as global hubs for communication, enabling individuals worldwide to exchange ideas, share experiences, and engage in discussions (Gruzd and Haythornthwaite, 2013). Similarly, in the realm of software development, platforms like GitHub have flourished, providing developers from diverse backgrounds with the opportunity to contribute code, discuss technical matters, and address software issues, regardless of their geographic location (Booch and Brown, 2003; Sengupta et al., 2006). As of now, GitHub hosts a vast array of projects spanning various domains, including operating systems, programming languages, and application software, forming a vast and diverse open-source ecosystem (Tsay et al., 2014). Moreover, with the rise of crowdsourced software development, an increasing number of projects leverage large-scale, decentralized, and heterogeneous resources and manpower for development (Lin, 2018; Wang et al., 2022). This trend further expands and deepens the scope of open collaboration.
In such development environments, software quality assurance faces both new challenges and opportunities. The distributed nature of open collaborative development, coupled with diverse contributors and rapid iterations, necessitates novel approaches to ensure the quality and reliability of software products. Additionally, the sheer volume and complexity of open-source projects demand sophisticated quality assurance techniques to effectively manage and mitigate potential risks and issues.
@@
There exist notable roadmaps that have significantly contributed to the understanding and advancement of software quality assurance practices. One such landmark roadmap is “Software Testing: A Research Travelogue (2000–2014),” spearheaded by Professor Alessandro Orso and Gregg Rothermel (Orso and Rothermel, 2014). This seminal work primarily explores software testing techniques and methodologies, offering valuable insights into the evolution of testing technologies over the past decade and a half. By comparison, our focus in this work is on quality assurance, and the factors influencing software quality extend far beyond testing techniques alone. They encompass a myriad of aspects, including process considerations, the dynamics of involved personnel, and more. Particularly, within open collaborative software development paradigms, traditional process techniques may find limited applicability, while human factors become increasingly pronounced. Moreover, while software has long claimed to be a socio-technical system (Mens, 2016), it is within open communities where the true complexity of this socio-technical system’s impact on software quality becomes strikingly apparent. Therefore, a comprehensive overview and roadmap to software quality assurance in open collaborative environments demands a holistic understanding that encompasses not only testing methodologies but also broader contextual factors.
@@
Figure 1. Overview about software quality assurance under open collaborative development environment
@@
Taken in this sense, our exploration into software quality assurance within open collaborative development environments is guided by the recognition that three key dimensions — process, personnel, and technology — form a triad of fundamental factors, influencing software quality. These dimensions represent orthogonal facets that collectively shape the landscape of software development and quality assurance, as visually demonstrated in Figure 1. Each dimension offers unique insights into the challenges and opportunities inherent in ensuring software quality within open collaborative environments. By exploring these three dimensions—process, personnel, and technology—we achieve a holistic grasp of the intricate landscape of software quality assurance within open collaborative development environments. Together, they constitute the foundation of present-day effective methodologies for upholding software quality.
@@
Additionally, we also outline future trends. As emerging technologies like pre-trained Large Language Models (LLMs) continue to evolve, the landscape of software quality assurance research is expected to expand to encompass various aspects. This may include leveraging LLMs for enhancing testing practices, exploring testing methodologies tailored for LLMs and related applications, ensuring quality assurance for auto-generated code produced by AI models, addressing security concerns within AI model ecosystems, and refining process management and collaboration strategies within AI model-centric paradigms. These areas represent promising avenues for future research and innovation in the field of software quality assurance within open collaborative development environments.
@@
The paper is structured as follows: Section 2 to 4 provide an overview of quality assurance researches from the perspectives of process, personnel, and technology, respectively. Section 5 explores the opportunities and challenges in this field, while Section 6 concludes the paper.
@@ 2. Process Related
Process, being the cornerstone of software development, encompasses methodologies, workflows, and practices that govern quality assurance in open collaborative environments. It delves into practices like continuous integration, aimed at frequent code integration and automated testing to uphold quality standards. However, challenges lie in optimizing regression testing and prioritizing test cases for timely developer feedback. Additionally, modern code review practices leverage automation and machine learning to streamline the review process and enhance overall code quality. Community involvement further shapes the process dimension, bringing multidimensional knowledge as well as a plethora of noise and redundant information, posing significant challenges to process management.
2.1. Continuous Integration and Modern Code Review
@@
Continuous Integration (CI) is a software development practice designed to encourage development teams to integrate their code frequently into a shared version control repository through automated build and testing processes. Each integration is automatically built and usually validated by regression testing, upon completion of which the developers are provided feedback. The execution of regression tests may require significant computational resources and take hours or even days to be completed. The prolonged execution of regression tests can delay the CI cycles and prevent timely feedback to the developers (Yaraghi et al., 2022). Test case selection and prioritization techniques have been proposed to improve regression testing by selecting and prioritizing test cases in order to provide early feedback to developers. We introduce the Test Case Prioritization (TCP) techniques which share similarities with selection techniques.
@@
Earlier attempts start with heuristics-based techniques, e.g., Elbaum et al. (Elbaum et al., 2014) prioritized tests based on whether they have not been executed for long or have failed in the recent commits, and Haghighatkhah et al. (Haghighatkhah et al., 2018) scheduled tests based on the combination of their previous execution results and similarity. Later, other researchers harness the power of machine learning by using a large amount of historical data in CI, and propose numerous machine learning based TCP techniques which have been demonstrated to be promising. They build neural models to predict the optimal sequence of tests instead of human-defined strategies. In particular, these machine learning based techniques can be categorized into supervised learning-based (Busjaeger and Xie, 2016; Bertolino et al., 2020; Sharif et al., 2021; Yaraghi et al., 2022), and reinforcement learning-based techniques (Spieker et al., 2017; Bagherzadeh et al., 2022; Lima and Vergilio, 2022). The later ones would continuously adjust its prioritization strategy, i.e., it is first tested (i.e., prioritizing the test suite) and then trained based on the prioritization feedback.
Continuous integration and modern code review are complementary practices that work together to ensure code quality, foster collaboration, and drive continuous improvement in software development projects. Modern code review involves systematically reviewing and evaluating proposed changes to the source code of a software project. Reviewers need to thoroughly understand the source code during code review activities and leave review comments to interact with developers, which involves a significant amount of effort and time of reviewers. Researchers are paying attention to automating code review activities, including comment location prediction (Hellendoorn et al., 2021; Shi et al., 2019), review comment recommendation (Gupta and Sundaresan, 2018; Li et al., 2022a, b) and code refinement (Tufano et al., 2021). Specifically, Shi et al. (Shi et al., 2019) proposed the DACE framework based on CNN and LSTM to predict whether a hunk in the code change will be accepted by the reviewers. Hellendoorn et al. (Hellendoorn et al., 2021) used the Transformer architecture to solve this task. To save the time reviewers spent on writing reviews related to common issues such as coding style and documentations, Gupta and Sundaresan (Gupta and Sundaresan, 2018) proposed the LSTM-based model DeepMem. It learns the relations between code changes and review comments, and recommends review comments automatically based on existing code reviews and code changes. CodeReviewer (Li et al., 2022a) and AUGER (Li et al., 2022b) utilize the pre-training models for the automatic generation of review comments. Tufano et al. (Tufano et al., 2021) used deep learning techniques to automate a different task in code review. They train a Transformer model to revise the contributors’ code to implement the requirements from the review comments.
2.2. Community Involvement and Information Filtering
Diverse individuals with different roles and identities come together to contribute towards open source software development. This collaborative effort encompasses developers, testers, designers, project managers, and end-users, among others, who collectively engage in activities such as issue reporting, code review, and discussions to drive the improvement and refinement of software products. However, as participation in these collaborative platforms grows, it inevitably leads to the generation of a vast amount of information, including issue reports, discussions, and contributions. While this influx of information can provide valuable insights and perspectives, it also poses challenges related to information overload, redundancy, and quality assurance. To address these challenges, various information filtering techniques have emerged within open collaborative development environments. These techniques aim to streamline the information flow, improve signal-to-noise ratio, and enhance the overall quality of contributions.
@@ 2.2. Community Involvement and Information Filtering
Duplicate detection is the most-commonly employed technique, which aims at identifying and eliminating redundant or duplicate information, such as duplicate issue reports or discussions on similar topics. For example, Nguyen et. al. (Nguyen et al., 2012) applied information retrieval techniques for duplicate detection by computing the textual similarity between two reports. Sun et. al. (Sun et al., 2010) designed a set of features for measuring the reports’ similarity in terms of textual descriptions and attributes, and employed machine learning techniques for duplicate detection. Yang et. al. (Yang et al., 2016) modeled the semantic similarity of reports using deep learning techniques for duplicate detection. In addition, Huang et. al. (Huang et al., 2020) and Zhang et. al. (Zhang et al., 2023a) respectively conducted experimental evaluations of the commonly-used approaches for duplicate detection.
Knowledge extraction from developer conversations also plays an increasingly significant role for accelerating the development process. Di Sorbo et al. (Sorbo et al., 2015) proposed a taxonomy of intentions to classify sentences in developer mailing lists. Qu et al. (Qu et al., 2019) utilized machine learning methods to perform user intent prediction. Chowdhury et al. (Chowdhury and Hindle, 2015) filtered out off-topic discussions in programming IRC channels by engaging Stack Overflow discussions as positive examples and YouTube video comments. Shi et al. (Shi et al., 2021) extracted the issue-solution pairs from massive chat messages to facilitate OSS developers to check and fix issues.
@@ 3. Personnel Related
In the realm of open-source software development, the involvement and collaboration of individuals play a crucial role in shaping the quality and success of projects. This dimension delves into various aspects concerning individuals, including code review practices, worker recommendation for issues and testing tasks, and support for newcomers.
3.1. Code Reviewer Recommendation
Modern code review involves two roles - code contributors and code reviewers. In a simplified code review procedure, code contributors commit code changes and propose a code review request; code reviewers respond to the request asynchronously, then review code changes and give feedback. However, practitioners often find reviewers manually or, at best, semi-manually in practice. Therefore, researchers proposed code reviewer recommendation approaches to recommend code reviewers efficiently. These approaches are achieved by learning the historical data of code review activity, including the developers’ code changes and participation in prior code reviewers (Chen et al., 2022; Hu et al., 2020; Rong et al., 2022). For example, Thongtanunam et al. (Thongtanunam et al., 2015) leverages file paths for expertise-based reviewer recommendation, which assumes that the files located in close files may share similar functionality and are likely to be reviewed by reviewers with common experience. Zanjani et al. (Zanjani et al., 2015) considers the reviewing history (review number, review time). It counts the number of comments to the file as part of scores. It also considers the frequency and recency of historical contributors to measure the reviewer expertise and give a final recommendation. Motivated by the fact that a revision may have multiple reviewers and they may impact each other through publicly posted comments, Rong et al. (Rong et al., 2022) adopted the hypergraph technique to model this high-order relationship and developed HGRec.
@@ 3.2. Issue Triaging
Various issues appear during software development and maintenance, and issue fixing is a time-consuming and costly task. Once a issue report is received, assigning it to a suitable developer within a short time interval can reduce the time and cost of the issue fixing process. This assignment process is known as issue triaging. Issue triaging is a time-consuming process since often a large number of developers are involved in software development and maintenance. To aid in finding appropriate developers, earlier practice adopted techniques as machine learning, graph analysis, fuzzy set, and topic modeling. For example, Anvik et al. (Anvik et al., 2006) utilized machine learning methods to solve it. Jeong et al. (Jeong et al., 2009) proposed to use a bug tossing graph to improve issue triaging prediction accuracy. Tamrawi et al. (Tamrawi et al., 2011) proposed a method called Bugzie, which uses a fuzzy set and cache-based approach to increase the accuracy of issue triaging. Xia et al. (Xia et al., 2017) proposed a specialized topic modeling algorithm named multi-feature topic model for issue triaging. Later studies used deep learning for bug triaging, e.g., Lee et al. (Lee et al., 2017) applied word embedding to train a CNN based classifier. Dipongkor et al. (Dipongkor and Moran, 2023) conducted the experimental evaluation about fine-tuning the transformer-based language models for this task.
3.3. Crowd Tester Recommendation
With the emergence of crowdsourced software development, more and more projects are utilizing extensive, decentralized, and diverse resources and manpower for their development efforts. Among them, a significant category is crowdsourced software testing (also short for crowdtesting). Unlike typical software development projects, testing tasks typically require a group of testers, ideally with diverse backgrounds, to achieve the goal of diversified testing and covering different areas of the software. Therefore, different from the worker recommendation for general crowdsourcing tasks, the tester recommendation for crowdtesting tasks tend to take into account the diversity of the recommended workers. For example, Wang et al. (Wang et al., 2021) propose a multi-objective crowd tester recommendation approach, which aims at recommending crowd tester by maximizing the bug detection probability of testers, the relevance with the test task, the diversity of testers, and minimizing the test cost. Additionally, previous studies on worker recommendation mainly focus on one-time recommendations with respect to the initial context at the beginning of a new task. However, for crowdtesting, a typical task can last from 3 days to 2 weeks, during when crowd testers can freely conduct the testing and submit reports. Taking this into account, Wang et al. (Wang et al., 2020) point out the need for accelerating crowdtesting by recommending appropriate testers in a dynamic manner, and they propose a context-aware in-process crowd testers recommendation approach, to detect more bugs earlier and potentially shorten the testing period.
3.4. Support for New Comers
A continuous influx of newcomers is essential for the survival, long-term success, and continuity of OSS projects [1, 8]. However, researches pointed out that the lack of appropriate tasks, up-to-date documents, new tools are the factors that might hinder developers from joining the project (Shibuya and Tamai, 2009; Zhou and Mockus, 2012). To help newcomers find their tasks, Huang et al. (Huang et al., 2021b) defined 79 features to characterize the suitable tasks, i.e., good first issue (GFIs), and built machine learning models to predict the GFIs with the proposed features. Similarly, Xiao et al. (Xiao et al., 2022) modeled an issue with features from multiple dimensions (content, background, and dynamics) and built the classifier to generate its probability of being a GFI. Following that, Tan et al. (Tan et al., 2023) investigated whether it is enough to recommend tasks to newcomers. They analyze the resolution process of GFIs, and find that expert involvement positively correlates with newcomers’ successful contributions. Canfora et al. (Canfora et al., 2012) focused on identifying and recommending mentors in software projects by mining data from mailing lists and versioning systems.

### v2

# A Roadmap for Software Testing in Open-Collaborative and AI-Powered Era

@@ Abstract.
Internet technology has given rise to an open-collaborative software development paradigm, necessitating the open-collaborative schema to software testing. It enables diverse and globally distributed contributions, but also presents significant challenges to efficient testing processes, coordination among personnel, and management of testing artifacts. At the same time, advancements in artificial intelligence (AI) have enhanced testing capabilities and enabling automation, while also introducing new testing needs and unique challenges for AI-based systems. In this context, this paper explores software testing in the open-collaborative and AI-powered era, focusing on the interrelated dimensions of process, personnel, and technology. Among them, process involves managing testing workflows and artifacts to improve efficiency, personnel emphasizes the role of individuals in ensuring testing quality through collaboration and contributions, while technology refers to AI methods that enhance testing capabilities and address challenges in AI-based systems. Furthermore, we delve into the challenges and opportunities arising from emerging technologies such as large language models (LLMs) and the AI model-centric development paradigm.
Software Testing, Artificial Intelligence, AI, Large Language Model, LLM, Open Source, Open Collaborative
$c$s: Software and its engineering Software creation and management
@@ 1. Introduction
Open collaboration is a hallmark of the Internet era in the past ten to twenty years, where geographical distances cease to be a barrier. Platforms like social media networks and online forums have emerged as global hubs for communication, enabling individuals worldwide to exchange ideas, share experiences, and engage in discussions (Gruzd and Haythornthwaite, 2013). Similarly, in the realm of software development, this has led to the rise of the open collaborative development paradigm (Crowston et al., 2008). Platforms like GitHub have flourished, providing developers from diverse backgrounds with the opportunity to contribute code, discuss technical matters, and address software issues, regardless of their geographic location (Booch and Brown, 2003; Sengupta et al., 2006). In such development environments, software testing and quality assurance faces both new challenges and opportunities. Unlike the closed-environment testing paradigm with fixed contributors and predefined setups, open-environment testing faces the challenges of distributed collaboration, diverse contributors, and rapid iterations, often resulting in uneven outcomes. This variability necessitates not only a focus on testing techniques but also on the testing process itself, such as improving coordination of testing activities to handle variability and streamlining the management of testing artifacts to achieve better cost-effectiveness. Additionally, attention must also be given to the personnel involved, providing them with the support needed to enhance their efficiency.
Another significant shift is the rise of artificial intelligence (AI) technologies. From the early days of machine learning and deep learning to the advent of pre-trained large language models (LLMs), these advancements offer effective solutions to address the challenges inherent in open-collaborative testing.
@@
There exist notable roadmaps that have significantly contributed to the understanding and advancement of software testing practices. One such landmark roadmap is “Software Testing: A Research Travelogue (2000–2014),” spearheaded by Professor Alessandro Orso and Gregg Rothermel (Orso and Rothermel, 2014). This seminal work primarily explores software testing techniques and methodologies, offering valuable insights into the evolution of testing technologies over the past decade and a half.
However, a decade has passed since then, and with the constant emergence of new technologies, there is a pressing need for a new roadmap to summarize and project future research in software testing. Moreover, we aim to adopt a broader perspective, recognizing that the factors influencing software testing extend well beyond the techniques alone. Therefore, a comprehensive overview and roadmap for software testing in the open-collaborative and AI-powered era requires a holistic understanding that encompasses not only testing methodologies but also broader contextual factors.
@@
Figure 1. Overview about software testing in open-collaborative and AI-powered era
@@
Taken in this sense, our exploration of software testing in the open-collaborative and AI-powered era is guided by the recognition that three key dimensions—process, personnel, and technology—form a triad of fundamental factors influencing software testing practices. These dimensions represent orthogonal facets that collectively shape the landscape of software testing, as visually demonstrated in Figure 1. Among them, the process dimension focuses on methodologies to manage dynamic, distributed contributions efficiently and ensure timely, comprehensive testing coverage. The personnel dimension highlights the critical role of human contributions in identifying and resolving issues to maintain testing quality. The technology dimension emphasizes the impact of AI advancements in enhancing testing capabilities and addressing the unique challenges of AI-driven systems. By examining these three dimensions—process, personnel, and technology—we gain a comprehensive understanding of the intricate landscape of software testing within open-collaborative environments.
@@
Additionally, we also outline future trends. As emerging technologies like LLMs continue to evolve, the landscape of software testing research is expected to expand to encompass various aspects. This may include leveraging LLMs for enhancing testing practices, exploring testing methodologies tailored for LLMs and related applications, ensuring quality assurance for auto-generated code produced by AI models, and refining collaboration strategies within AI model-centric paradigms. These areas represent promising avenues for future research and innovation in the field of software testing under open collaborative and AI-powered era.
To curate the relevant works, we adopt an informal methodology informed by the authors’ expertise in software testing and AI. The selection prioritized survey papers and high-impact studies from top-ranked venues in software engineering as well as AI related venues. This allowed us to quickly incorporate emerging insights into our discussion while focusing on the core challenges and opportunities related to software testing in the AI era.
@@
The paper is structured as follows: Section 2 to 4 provide an overview of software testing researches from the perspectives of process, personnel, and technology, respectively. Section 5 explores the opportunities and challenges in this field, while Section 6 concludes the paper.
@@ 2. Process Related
In open-collaborative environments, the process dimension of software testing involves practices and methodologies that enable efficient and effective testing in dynamic, distributed development settings, alongside the streamlined management of testing artifacts.
2.1. Testing Process Management
@@
In modern software development, Continuous Integration (CI) has become a cornerstone practice for maintaining code quality and fostering collaboration. By frequently integrating code into a shared version control repository, CI ensures that each change is automatically validated through automated build and testing pipelines. A critical aspect of this process is regression testing, which serves as a key manifestation of the testing process by verifying that new changes do not introduce defects into existing functionality. While regression testing is invaluable for maintaining software quality, executing a full suite of tests can be highly resource-intensive and time-consuming, often requiring hours or even days to complete. This delay can hinder the CI cycle and prevent timely feedback for developers (Yaraghi et al., 2022). To mitigate these challenges, techniques such as test case prioritization and test case selection have been developed to optimize the testing process, by selecting and prioritizing test cases in order to provide early feedback to developers. We introduce the Test Case Prioritization (TCP) techniques which share similarities with selection techniques.
@@
Earlier attempts start with heuristics-based techniques, e.g., Elbaum et al. (Elbaum et al., 2014) prioritized tests based on whether they have not been executed for long or have failed in the recent commits, and Haghighatkhah et al. (Haghighatkhah et al., 2018) scheduled tests based on the combination of their previous execution results and similarity. Later, other researchers harness the power of machine learning by using a large amount of historical data in CI, and propose numerous machine learning based TCP techniques which have been demonstrated to be promising. They build neural models to predict the optimal sequence of tests instead of human-defined strategies. In particular, these machine learning based techniques can be categorized into supervised learning-based (Busjaeger and Xie, 2016; Bertolino et al., 2020; Sharif et al., 2021; Yaraghi et al., 2022), and reinforcement learning-based techniques (Spieker et al., 2017; Bagherzadeh et al., 2022; Lima and Vergilio, 2022). The latter ones would continuously adjust its prioritization strategy, i.e., it is first tested (i.e., prioritizing the test suite) and then trained based on the prioritization feedback.
Apart from that, in open-collaborative environments, crowdsourced testing (also typically abbreviated as crowdtesting) has become a key strategy for enhancing software quality by leveraging contributions from a large, diverse group of external testers. Crowdtesting is hard to manage in nature. Given the unpredictability of distributed crowdtesting processes, it is difficult to estimate (a) remaining number of bugs yet to be detected or (b) required cost to find those bugs. Experience-based decisions may result in ineffective crowdtesting processes, e.g., there is an average of 32% wasteful spending in current crowdtesting practices. Wang et. al. (Wang et al., 2019) explored automated decision support to effectively manage crowdtesting processes.
2.2. Testing Artifacts Management
@@
In open-collaborative environments, the testers—often from different backgrounds and geographical locations—can provide valuable feedback by identifying issues that might not be caught through traditional testing methods. However, the volume of contributions from such a wide range of sources presents unique challenges. The influx of issue reports, bug findings, and feedback (all of them can be treated as testing artifacts) can lead to information overload, where valuable insights are buried under redundant or conflicting contributions. This makes it difficult to prioritize and filter out irrelevant data, which can undermine the efficiency of the testing process. Effective information filtering techniques are therefore necessary to improve the signal-to-noise ratio and streamline the testing process. Duplicate detection is the most-commonly employed technique, which aims at identifying and eliminating redundant or duplicate information, such as duplicate issue reports or discussions on similar topics. For example, Nguyen et. al. (Nguyen et al., 2012) applied information retrieval techniques for duplicate detection by computing the textual similarity between two reports. Sun et. al. (Sun et al., 2010) designed a set of features for measuring the reports’ similarity in terms of textual descriptions and attributes, and employed machine learning techniques for duplicate detection. Yang et. al. (Yang et al., 2016) modeled the semantic similarity of reports using deep learning techniques for duplicate detection. In addition, Huang et. al. (Huang et al., 2020) and Zhang et. al. (Zhang et al., 2023) respectively conducted experimental evaluations of the commonly-used approaches for duplicate detection.
@@ 3. Personnel Related
In open-collaborative software testing, the involvement of individuals plays a crucial role in the success of the testing process. Human factors influence the quality of testing, as contributions and collaborations from testers, developers, and stakeholders are vital for identifying critical issues, quickly fix the issues, and ensuring comprehensive coverage.
3.1. Tester Recommendation
Unlike typical software development projects, testing tasks typically require a group of testers, ideally with diverse backgrounds, to achieve the goal of diversified testing and covering different areas of the software. Therefore, different from the worker recommendation for general crowdsourcing tasks, the tester recommendation for crowdtesting tasks tend to take into account the diversity of the recommended workers. For example, Wang et al. (Wang et al., 2021) propose a multi-objective crowd tester recommendation approach, which aims at recommending crowd tester by maximizing the bug detection probability of testers, the relevance with the test task, the diversity of testers, and minimizing the test cost. Additionally, previous studies on worker recommendation mainly focus on one-time recommendations with respect to the initial context at the beginning of a new task. However, for crowdtesting, a typical task can last from 3 days to 2 weeks, during which crowd testers can freely conduct the testing and submit reports. Taking this into account, Wang et al. (Wang et al., 2020) point out the need for accelerating crowdtesting by recommending appropriate testers in a dynamic manner, and they propose a context-aware in-process crowd testers recommendation approach, to detect more bugs earlier and potentially shorten the testing period.
@@ 3.2. Issue Triaging
Various issues appear during software testing, and issue fixing is a time-consuming and costly task. Once an issue report is received, assigning it to a suitable developer within a short time interval can reduce the time and cost of the issue fixing process. This assignment process is known as issue triaging. Issue triaging is a time-consuming process since often a large number of developers are involved in software testing. To aid in finding appropriate developers, earlier practice adopted techniques as machine learning, graph analysis, fuzzy set, and topic modeling. For example, Anvik et al. (Anvik et al., 2006) utilized machine learning methods to solve it. Jeong et al. (Jeong et al., 2009) proposed to use a bug tossing graph to improve issue triaging prediction accuracy. Tamrawi et al. (Tamrawi et al., 2011) proposed a method called Bugzie, which uses a fuzzy set and cache-based approach to increase the accuracy of issue triaging. Xia et al. (Xia et al., 2017) proposed a specialized topic modeling algorithm named multi-feature topic model for issue triaging. Later studies used deep learning for bug triaging, e.g., Lee et al. (Lee et al., 2017) applied word embedding to train a CNN based classifier. Dipongkor et al. (Dipongkor and Moran, 2023) conducted the experimental evaluation about fine-tuning the transformer-based language models for this task.
3.3. Human–computer Collaborative Testing
There are studies that analyze incorporating automation technologies to assist manual testing, exemplifying human-computer interaction in the software testing process. For instance, Liu et al. automatically trace testers’ actions and use explicit visual annotations to guide or remind them of unexplored areas, helping testers avoid missing functionalities or repeating steps (Liu et al., 2022). Similarly, Chen et al. investigate human collaboration in crowd testing scenarios. They utilize interactive event-flow graphs to track and aggregate each tester’s interactions into a single directed graph, which visualizes the test cases already explored. Crowd testers can interact with these graphs to discover new navigation paths and improve test coverage (Chen et al., 2020).